OpenStudy (dls):

Prove that.. @yrelhan4 @Mashy

OpenStudy (dls):

\[\LARGE U=\sqrt{2g(H+\frac{R^2}{16H})}\]

OpenStudy (dls):

u=projectile wala..

OpenStudy (anonymous):

\[H = \frac{u^2\sin^2( \theta)} {2g}... R = \frac{u^2\sin(2 \theta)}{g}\]

OpenStudy (dls):

thanks! :D

OpenStudy (dls):

@Mashy PLEASE GIVE MY EXAM TOMORROW!! DX

OpenStudy (dls):

??

OpenStudy (anonymous):

ok ;p